OpenStudy (anonymous):

I'm having a problem with solving this double integral...can someone help me?

OpenStudy (anonymous):

Hard to tell if we can't see it ;)

OpenStudy (anonymous):

ok there's no way i can put it up like im seeing it because there's integral sign but i'll try my best...

OpenStudy (anonymous):

integral 0-2 (integral 0-x/2 5(e)^(x)^2 dy dx)...help me please

myininaya (myininaya):

first we integrate inside ignore the outside for right now

myininaya (myininaya):

since we are integrating with repsect to y we treat everything thats not y ish as a constant

OpenStudy (anonymous):

i under stand that. There's no y so I dont understand if i just put a y in there?

myininaya (myininaya):

what? so you are saying you don't see the inside being 5ye^x^2?

OpenStudy (anonymous):

no, there's not a y in the whole problem

myininaya (myininaya):

whats the antiderivative of 5 with repect to x

OpenStudy (anonymous):

0

myininaya (myininaya):

no

OpenStudy (anonymous):

5x?

myininaya (myininaya):

its 5x since the derivative of 5x is 5

myininaya (myininaya):

so now you understand how we got 5ye^x^2 for inner integral (we haven't used upper and lower limit yet for inner integral)

OpenStudy (anonymous):

ok so you added a y because we were doing respect to y?

myininaya (myininaya):

yes and the derivative of 5ye^x^2 with respect to y will be 5e^x^2

myininaya (myininaya):

so the inside integral becomes (5(x/2)e^x^2-0)

myininaya (myininaya):

now we have the outside integral that we integrating with respect to x

myininaya (myininaya):

use a substitution for x^2

OpenStudy (anonymous):

du/2=x

myininaya (myininaya):

du/2=x dx yes

myininaya (myininaya):

dont forget to put the limits in u terms

myininaya (myininaya):

so the upper limit will be 4 and the lower limit will be 0

myininaya (myininaya):

so we have int((5/4) *e^u, u=0..4)=(5/4)e^u, u=0..4= (5/4)e^4-(5/4)e^0=(5/4)e^4

myininaya (myininaya):

-1

myininaya (myininaya):

i forgot about the minus 1 sorry

myininaya (myininaya):

e^0=1

OpenStudy (anonymous):

thank you very much

myininaya (myininaya):

np don't be afraid to ask any other questions or this one over if you want

OpenStudy (anonymous):

is this a 24 hour website?

OpenStudy (anonymous):

on this problem, and when doing substitution is it always required to put the limits in u terms?

myininaya (myininaya):

i don't know. people come and go as they please. the tutors work whenever they want because we don't work for the website we work because we think its fun lol

myininaya (myininaya):

if you don't put the limits in terms of u, you have to change it back to x and use the limits for x

OpenStudy (anonymous):

is there any way you can explain that?

myininaya (myininaya):

its the same thing

myininaya (myininaya):

lets do an example...

OpenStudy (anonymous):

this seems like it's going to be a really hard problem

myininaya (myininaya):

no its not hard i'm just busy doing something else at sametime

myininaya (myininaya):

pretend we have int(5x,x=0..1)

myininaya (myininaya):

if we integrate this normal way we get 5x^2/2 x=0..1 = 5/2

myininaya (myininaya):

we could also do u substitution if we just wanted to for fun so let u=5x

myininaya (myininaya):

du=5dx => du/5=dx

OpenStudy (anonymous):

ok

myininaya (myininaya):

so we get int(1/5 (u), u=blah blah) =1/5 *u^2/2 let's change this back to x so we will have 1/5*((5x))^2/2 x=0..1 but we could have already done the 5x if we had used u=5x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it ok to ask another question or should i just retype it on the "ask a question"? you seem to really know your stuff thats why im asking

myininaya (myininaya):

that didn't answer it?

OpenStudy (anonymous):

i mean i have another question.

myininaya (myininaya):

oh lol yeah I have to do something. I'm fixing to have to move a mattress

OpenStudy (anonymous):

ok, thanks for all your help. hope you make lots of money with your math skill one day